Solve systems of three variables. linear eqations
2x+2y+z=1
-x+y+2z=3
x+2y+4z=0

Answer by TimothyLamb(4379)

---
2x+2y+z=1
-x+y+2z=3
x+2y+4z=0
---
copy and paste the above standard form linear equations in to this (matrix method) solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= -2
y= 3
z= -1
---
NOTE TO STUDENT:
now solve by elimination:
---
-x+y+2z=3
x+2y+4z=0
---
add this sub-system and call it A:
---
3y+6z=3
---
2x+2y+z=1
2(-x+y+2z=3)
---
add this sub-system and call it B:
---
4y+5z=7
---
the A and B sub-system:
---
3y+6z=3
4y+5z=7
---
add the A and B sub-system:
---
-4(3y+6z=3)
3(4y+5z=7)
---
-9z=9
---
z = -1
---
substitute z in B:
---
4y+5(-1)=7
4y=12
---
y = 3
---
substitute y and z in any of the original equations:
---
x+2(3)+4(-1)=0
x+6-4=0
---
x = -2
---
answer by elimination:
x= -2
y= 3
z= -1
NOTE: the matrix method (used by the above solver), and the elimination method agree
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php